### v3

# Improve the Precision of Area Under the Curve Estimation for Recurrent Events Through Covariate Adjustment

Abstract
The area under the curve (AUC) of the mean cumulative function (MCF) has recently been introduced as a novel estimand for evaluating treatment effects in recurrent event settings, offering an alternative to the commonly used Lin-Wei-Yang-Ying (LWYY) model. The AUC of the MCF provides a clinically interpretable summary measure that captures the overall burden of disease progression, regardless of whether the proportionality assumption holds. To improve the precision of the AUC estimation while preserving its unconditional interpretability, we propose a nonparametric covariate adjustment approach. This approach guarantees efficiency gain compared to unadjusted analysis, as demonstrated by theoretical asymptotic distributions, and is universally applicable to various randomization schemes, including both simple and covariate-adaptive designs. Extensive simulations across different scenarios further support its advantage in increasing statistical power. Our findings highlight the importance of covariate adjustment for the analysis of AUC in recurrent event settings, offering practical guidance for its application in randomized clinical trials.
Keywords Area Under the Curve  $\cdot$ Lin-Wei-Yang-Ying (LWYY) Model  $\cdot$ Mean Cumulative Function  $\cdot$ Nonparametric Covariate Adjustment  $\cdot$ Recurrent Events
1 Introduction
In recent years, recurrent event analysis has gained traction in clinical studies, particularly in chronic conditions such as multiple hospitalizations for heart failure, recurrent relapses in multiple sclerosis, or frequent exacerbation in asthma [Solomon et al., 2024, Rogers et al., 2014, Cohen and Rudick, 2011, Keene et al., 2007]. A widely adopted approach for analyzing recurrent event data in clinical trials is the Lin-Wei-Yang-Ying (LWYY) model [Lin et al., 2000]. The LWYY model assumes proportionality between the mean cumulative functions (MCFs) of two arms, implying a constant rate ratio over time. This allows treatment effects to be evaluated through the ratio of the MCFs. In practice, this assumption might be violated, particularly in complex scenarios where treatment effects are time-varying. For example, the recent FINEARTS-HF study demonstrated time-dependent variation in rate ratio estimates [Vaduganathan et al., 2024]. While the LWYY model offers some robustness to model misspecification [Lin and Wei, 1989]—and its constant estimate can be interpreted as a weighted average of time-varying effects—this weighted average may lack clear clinical interpretability in some contexts, in which more flexible estimands that do not assume proportionality may be preferable.
The recent ICH E9(R1) addendum also emphasizes the importance of a robust definition quantifying the treatment effects in the settings of recurrent events with minimal assumptions. In this context, the area under the curve (AUC)-based approach for recurrent events proposed by Claggett et al. [2022] has gained considerable attention, offering a clinically meaningful summary of the cumulative burden of recurrent events without assuming proportionality [Gregson et al., 2023]. By calculating the area under the MCF for each treatment arm, one can derive a summary measure that reflects the overall mean burden of events experienced by patients over the follow-up period. A larger AUC signifies a greater cumulative burden, and treatment comparisons based on the difference or ratio of AUCs allow for intuitive interpretation in terms of absolute or relative reductions in disease burden. AUCs can be consistently estimated by the Ghosh-Lin estimator irrespective of whether the proportionality assumption holds, making it a robust and interpretable estimand for analyzing recurrent event data [Ghosh and Lin, 2000].
Incorporating prognostic baseline covariates into the design and analysis of clinical trial data can improve the efficiency of data utilization, leading to more precise estimation and demonstration of treatment effects [Food and Drug Administration (2023), FDA, European Medicines Agency (2015), EMA]. There are several statistical considerations for better practice in covariate adjustment. First, covariate adjustment should enhance the precision of treatment effect estimation without altering the definition of the estimand. Furthermore, Food and Drug Administration (2023) [FDA] emphasized that the method used should allow for valid inference under approximately the same minimal statistical assumptions as those required for unadjusted estimation in a randomized trial, thereby avoiding reliance on additional assumptions. One potential approach to covariate adjustment for the AUC is a regression-based method. This approach shares a similar model structure with the LWYY model but replaces the rate function in the LWYY model with the AUC. However, the regression-based approach has notable limitations. First, when covariates are included in the model, the resulting estimates become non-collapsible [Hauck et al., 1998], targeting conditional rather than unconditional treatment effects, which might alter the estimand. Additionally, regression models impose restrictive assumptions, including proportionality, correct specification of covariate functional forms, and the exponential link function. Any of these assumptions may be violated in practice. If the regression model is misspecified, the efficiency gains from covariate adjustment cannot be guaranteed [Kong and Slud, 1997].
To address these challenges, we propose a framework for covariate adjustment in the AUC-based summary for recurrent events, extending the methods developed by Ye et al. [2022, 2023, 2024]. The primary objective is to enhance the precision of AUC-based summaries while addressing the limitations associated with the regression-based approach and other semiparametric covariate adjustment approaches. Importantly, the covariate-adjusted AUC-based summary preserves the estimand, targeting the unconditional treatment effect. The proposed approach is robust to model misspecification, ensuring that a consistent and valid estimator is obtained regardless of model correctness or the proportionality assumption. This approach guarantees asymptotic efficiency gains over the unadjusted method. Additionally, the approach is universally applicable to simple and covariate-adaptive randomization schemes, achieving optimal efficiency gains, unlike other covariate adjustment approaches that are only applicable to simple randomization [Lu and Tsiatis, 2008].
The remainder of the paper is organized as follows. Section 2 covers the covariate adjustment approach for the AUC estimand for recurrent events. Section 3 presents simulation studies that demonstrate the advantage of the covariate-adjusted AUC-based summary in increasing statistical power, and Section 4 applies the method to real data examples. Section 5 concludes with a discussion and future directions.
2 Area Under the Mean Cumulative Function
The area under the MCF was introduced [Claggett et al., 2022] to address the question: What is the cumulative burden of recurrent events over a fixed follow-up period? It quantifies this burden by capturing both the frequency and timing of recurrent events (e.g., repeated hospitalizations). Conceptually, each occurrence of a recurrent event increases the total time a subject spends in the undesirable condition, contributing to their cumulative burden, which is measured by event-time. The size of this contribution depends on timing: the earlier the event occurs, the more event-time it adds. The MCF tracks the average accumulation of such events across the study population. Therefore, the area under the MCF summarizes the mean total burden due to recurrent events, with a higher AUC indicating greater overall burden. This quantity provides a robust, model-free method for comparing treatment groups in terms of cumulative burden, with treatment effects expressed through differences or ratios of AUCs to reflect absolute or relative reductions in recurrent event burden. In what follows, we formally define the AUC estimand and describe how it can be estimated from censored recurrent event data.
2.1 Data and estimand
Let $D$ represent the time to a terminal event and write $N^{D}(t)=\mathbb{I}(D\leq t)$ as the counting process for $D$. Let $N^{*}(t)$ be the number of recurrent events up to time $t$. Let $C$ denote the independent censoring time. An observation consists of $\{N(\cdot),T,\delta\}$, where $N(t)=N^{*}(t\wedge C)$, $T=C\wedge D$, $\delta=\mathbb{I}(D\leq C)$, and $a\wedge b=\hbox{min}(a,b)$. The observed data set $\mathcal{D}=\{N_{i}(\cdot),T_{i},\delta_{i}\}$, for $i=1,\dots,n$, consists of independent realizations of $\{N(\cdot),T,\delta\}$.
Let $S(u)=\mathbb{P}(D\geq u)$ be the survival function, and $dR(u)=\mathbb{E}\{dN^{*}(u)\mid T\geq u\}=\mathbb{E}\{dN^{*}(u)\mid D\geq u\}$, where $dN^{*}(t)=\lim_{\Delta t\to 0}\{N^{*}(t+\Delta t^{-})-N^{*}(t^{-})\}$. Then, the mean cumulative function (MCF) is defined as
[…]
The area under the MCF over $[0,\tau]$ is
[…]
The estimand $U(\tau)$ can be interpreted as the mean total event-free time lost from recurrent events over the interval $[0,\tau]$, which captures the cumulative burden of recurrent events in a clinically interpretable way. $U(\tau)$ is a natural extension of the restricted mean survival time (RMST) from time-to-first event settings to the recurrent event setting [Kim et al., 2017]. Specifically, in the time-to-first event setting, the MCF can be considered as the expected value of the counting process for the death event at time $t$, which is equivalent to the probability of experiencing death by time $t$—that is, 1 minus the survival probability at time $t$. As a result, in the time-to-first event setting, the area under the MCF over the interval $[0,\tau]$ equals $\tau$ minus the RMST over $[0,\tau]$.
Denote $U_{0}(\tau)$ and $U_{1}(\tau)$ as the area under the MCF for placebo and treatment arms. The treatment effect can be quantified as either the difference, $\Delta U(\tau)=U_{1}(\tau)-U_{0}(\tau)$, or the ratio, $U_{1}(\tau)/U_{0}(\tau)$. For convenience in later derivations, we define $R(\tau)=\hbox{log}\frac{U_{1}(\tau)}{U_{0}(\tau)}$, which represents the log-transformed ratio of the areas under the MCF. Similar to how the RMST is estimated based on the Kaplan-Meier estimator, the AUC under each treatment group can be nonparametrically estimated by plugging in the Ghosh-Lin estimator [Ghosh and Lin, 2000]. An open-source implementation of this method is available at https://github.com/zrmacc/MCC.
An estimand specifies five key attributes: the treatment (the specific intervention being evaluated), population (the group of subjects to which the treatment effect applies), variable (the outcome or endpoint being measured), intercurrent events (events occurring after randomization that may affect the interpretation or occurrence of the outcome), and the population-level summary (how the variable is summarized to compare treatment conditions). The five attributes of our estimand are defined in Table LABEL:tab:estimand.
2.2 Nonparametric Covariate Adjustment
We extend the nonparametric covariate adjustment approach introduced by Ye et al. [2022, 2023, 2024] to the AUC-based summary. The key concept involves linearizing the test statistic to derive an outcome for each patient, followed by the application of generalized regression adjustment or augmentation to these derived outcomes [Zhang et al., 2008, Tsiatis et al., 2008]. In this section, we focus on covariate adjustment for the ratio of AUCs. The corresponding method for the difference of AUCs follows analogously and is provided in Section A of the Supplementary Materials.
2.2.1 Preparation
Let $\hat{\mu}(t)$ and $\hat{U}(\tau)$ denote estimates of $\mu(t)$ and $U(\tau)$, obtained by substituting $S(u)$ with the Kaplan-Meier estimator $\hat{S}(u)$ and $dR(u)$ with the Nelson-Aalen estimator $d\hat{R}(u)=\sum_{i=1}^{n}dN_{i}(u)\big{/}\sum_{i=1}^{n}\mathbb{I}(T_{i}\geq u)$, where $N_{i}(t)=N_{i}^{*}(t\wedge C_{i})$ and $dN_{i}(t)=\lim_{\Delta t\to 0}\{N_{i}(t+\Delta t^{-})-N_{i}(t^{-})\}$.
[…]
$\sqrt{n}\{\hat{U}(\tau)-U(\tau)\}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi_{i}(\tau)+o_{p}(1)$
[…]
The estimates $\hat{\psi}_{i}(\tau)$, $\hat{P}_{i}$, and $\hat{Q}_{i}$ are obtained by replacing $S(u)$ with $\hat{S}(u)$, $\mu(s)$ with $\hat{\mu}(s)$, and estimating $\mathbb{P}(T\geq u)$ as $\sum_{i=1}^{n}\mathbb{I}(T_{i}\geq u)/n$. The martingales are estimated by: $d\hat{M}_{i}(u)=dN_{i}(u)-\mathbb{I}(T_{i}\geq u)d\hat{R}(u)$, and $d\hat{M}_{i}^{D}(u)=dN_{i}^{D}(u)-\mathbb{I}(T_{i}\geq u)d\hat{A}^{D}(u)$, with $d\hat{A}^{D}(u)=\sum_{i=1}^{n}dN^{D}_{i}(u)\big{/}\sum_{i=1}^{n}\mathbb{I}(T_{i}\geq u)$.
2.2.2 Covariate-Adjusted AUCs
Let $\hat{R}(\tau)=\hbox{log}\frac{\hat{U}_{1}(\tau)}{\hat{U}_{0}(\tau)}$ and $\hat{\mathcal{U}}_{L,R}=\frac{n_{0}n_{1}}{n}\left\{\hat{R}(\tau)-R(\tau)\right\}$, where $n_{0}$ and $n_{1}$ represent the sample sizes for the placebo and treatment arms, respectively. The linearization of $\hat{\mathcal{U}}_{L,R}$ is as follows:
$\hat{\mathcal{U}}_{L,R}=\frac{1}{n}\sum_{i=1}^{n}\left\{I_{i}(\hat{P}^{R}_{i}-\hat{Q}^{R}_{i})-(1-I_{i})(\hat{P}^{R}_{i}-\hat{Q}^{R}_{i})\right\}+o_{p}(1)$
where $I_{i}=1$ if subject $i$ is in the treatment arm and 0 otherwise, $\hat{P}^{R}_{i}=\left\{\frac{n_{0}}{\hat{U}_{1}(\tau)}\right\}^{I_{i}}\left\{\frac{n_{1}}{\hat{U}_{0}(\tau)}\right\}^{1-I_{i}}\hat{P}_{i}$ and $\hat{Q}^{R}_{i}=\left\{\frac{n_{0}}{\hat{U}_{1}(\tau)}\right\}^{I_{i}}\left\{\frac{n_{1}}{\hat{U}_{0}(\tau)}\right\}^{1-I_{i}}\hat{Q}_{i}$. The unadjusted test statistic on the ratio of AUCs is $\mathcal{T}_{L,R}=\sqrt{n}\hat{\mathcal{U}}_{L,R}/\hat{\sigma}_{L,R}$, where
$\hat{\sigma}_{L,R}^{2}=\frac{1}{n}\left[\frac{n_{0}^{2}n_{1}}{\hat{U}_{1}(\tau)^{2}}\mathbb{E}\left\{\hat{\psi}_{1}(\tau)^{T}\hat{\psi}_{1}(\tau)\right\}+\frac{n_{1}^{2}n_{0}}{\hat{U}_{0}(\tau)^{2}}\mathbb{E}\left\{\hat{\psi}_{0}(\tau)^{T}\hat{\psi}_{0}(\tau)\right\}\right].$
Here, $\hat{\psi}_{1}(\tau)$ and $\hat{\psi}_{0}(\tau)$ consist of $\hat{\psi}_{i}(\tau)$ from subjects in the treatment and placebo arms, respectively. Derivations of $\hat{\mathcal{U}}_{L,R}$ and the asymptotic properties of $\mathcal{T}_{L,R}$ are provided in detail in Section B of the Supplementary Materials.
Let $\boldsymbol{X}^{*}$ be a $p$-dimensional vector of baseline covariates. Define $X_{i}=X_{i}^{*}-\bar{X}$, where $X_{i}^{*}$ is the $i$th column of $\boldsymbol{X}^{*}$, and $\bar{X}$ is the sample mean of all the $X_{i}^{*}$, ensuring $\mathbb{E}(X_{i})=0$. Following Ye et al. [2024], we treat $\hat{P}^{R}_{i}$ and $\hat{Q}^{R}_{i}$ as outcomes and apply the generalized regression adjustment or augmentation [Tsiatis et al., 2008]. This results in the covariate-adjusted $\hat{\mathcal{U}}_{CL,R}$:
$\hat{\mathcal{U}}_{CL,R}=\frac{1}{n}\sum_{i=1}^{n}\Bigg{[}I_{i}\left\{(\hat{P}^{R}_{i}-X_{i}^{T}\hat{\beta}_{1,P}^{R})-(\hat{Q}^{R}_{i}-X_{i}^{T}\hat{\beta}_{1,Q}^{R})\right\}-(1-I_{i})\left\{(\hat{P}^{R}_{i}-X_{i}^{T}\hat{\beta}_{0,P}^{R})-(\hat{Q}^{R}_{i}-X_{i}^{T}\hat{\beta}_{0,Q}^{R})\right\}\Bigg{]}=\hat{\mathcal{U}}_{L,R}-\hat{\mathcal{A}_{R}}$
where
$\hat{\mathcal{A}_{R}}$ $=$ $\frac{1}{n}\sum_{i=1}^{n}\left\{I_{i}X_{i}^{T}\left(\hat{\beta}_{1,P}^{R}-\hat{\beta}_{1,Q}^{R}\right)\right.-\left.(1-I_{i})X_{i}^{T}\left(\hat{\beta}_{0,P}^{R}-\hat{\beta}_{0,Q}^{R}\right)\right\},$
$\hat{\beta}_{j,P}^{R}$ $=$ $\left(\sum_{i:I_{i}=j}X_{i}X_{i}^{T}\right)^{-1}\sum_{i:I_{i}=j}X_{i}\hat{P}^{R}_{i},$
$\hat{\beta}_{j,Q}^{R}$ $=$ $\left(\sum_{i:I_{i}=j}X_{i}X_{i}^{T}\right)^{-1}\sum_{i:I_{i}=j}X_{i}\hat{Q}^{R}_{i}.$
Let $\hat{R}^{(a)}(\tau)$ denote the adjusted $\hat{R}(\tau)$, and is given by $\hat{R}^{(a)}(\tau)=\hat{R}(\tau)-\frac{n}{n_{0}n_{1}}\hat{\mathcal{A}_{R}}$. Then, $\hat{\mathcal{U}}_{CL,R}=\frac{n_{0}n_{1}}{n}\left\{\hat{R}^{(a)}(\tau)-R(\tau)\right\}$. Define the covariate-adjusted test statistic on the ratio of AUCs as $\mathcal{T}_{CL,R}=\sqrt{n}\hat{\mathcal{U}}_{CL,R}/\hat{\sigma}_{CL,R}$, where $\hat{\sigma}_{CL,R}^{2}=\hat{\sigma}_{L,R}^{2}-\frac{n_{0}n_{1}}{n^{2}}(\hat{\beta}_{1,P}^{R}-\hat{\beta}_{1,Q}^{R}+\hat{\beta}_{0,P}^{R}-\hat{\beta}_{0,Q}^{R})^{T}\sum_{X}(\hat{\beta}_{1,P}^{R}-\hat{\beta}_{1,Q}^{R}+\hat{\beta}_{0,P}^{R}-\hat{\beta}_{0,Q}^{R})$. The theorem below establishes the asymptotic properties of $\hat{\mathcal{U}}_{CL,R}$ and $\mathcal{T}_{CL,R}$. The technical proofs are included in Section C of the Supplementary Materials.
Theorem 1
The following results hold under both simple randomization and covariate-adaptive randomization schemes:
Under the $H_{0,R}:R(\tau)=0$ or the alternative hypothesis, $\sqrt{n}\hat{\mathcal{U}}_{CL,R}\stackrel{{\scriptstyle d}}{{\to}}N\left(0,\sigma_{CL,R}^{2}\right)$, $\hat{\sigma}_{CL,R}^{2}\stackrel{{\scriptstyle p}}{{\to}}\sigma_{CL,R}^{2}$, and $\mathcal{T}_{CL,R}\stackrel{{\scriptstyle d}}{{\to}}N(0,1)$, where $\sigma_{CL,R}^{2}=\sigma_{L,R}^{2}-\frac{n_{0}n_{1}}{n^{2}}(\beta_{1,P}^{R}-\beta_{1,Q}^{R}+\beta_{0,P}^{R}-\beta_{0,Q}^{R})^{T}\sum_{X}(\beta_{1,P}^{R}-\beta_{1,Q}^{R}+\beta_{0,P}^{R}-\beta_{0,Q}^{R})$.
$H_{0,R}$ is rejected when $|\mathcal{T}_{CL,R}|>z_{\alpha/2}$ and the adjusted $100(1-\alpha)\%$ CI for $R(\tau)$ is $\big{(}\hbox{exp}\big{[}\hat{R}^{(a)}(\tau)-z_{\alpha/2}SE\left\{\hat{R}^{(a)}(\tau)\right\}\big{]},\hbox{exp}\big{[}\hat{R}^{(a)}(\tau)+z_{\alpha/2}SE\left\{\hat{R}^{(a)}(\tau)\right\}\big{]}\big{)}$, where $SE\left\{\hat{R}^{(a)}(\tau)\right\}=\left(\frac{n}{n_{0}^{2}n_{1}^{2}}\hat{\sigma}_{CL,R}^{2}\right)^{1/2}$.
As proved by Ye et al. [2024], $\sigma_{CL,R}^{2}\leq\sigma_{L,R}^{2}$, with strict inequality holding unless either (i) the covariates $X_{i}^{*}$ are uncorrelated with $P_{i}^{R}$ and $Q_{i}^{R}$, or (ii) the covariates not used in the randomization are uncorrelated with $P_{i}^{R}$ and $Q_{i}^{R}$, given the covariates used in the randomization. Therefore, the adjusted $\mathcal{T}_{CL,R}$ provides a guaranteed efficiency gain over the unadjusted $\mathcal{T}_{L,R}$.
3 Simulations
3.1 Data Generation
Three independent baseline covariates $\boldsymbol{X}^{*}=(X_{1}^{*},X_{2}^{*},X_{3}^{*})$ are considered. Specifically, $X_{1}^{*}$ follows a Bernoulli distribution with probability of 0.5 while $X_{2}^{*}$ and $X_{3}^{*}$ independently follow a normal distribution with mean of 0 and standard deviation of 2.
Two randomization schemes with equal allocation are considered. The first scheme is simple randomization, where treatments are assigned to subjects completely at random. This approach may result in imbalances across the treatment arms concerning $(X_{1}^{*},X_{2}^{*},X_{3}^{*})$. To mitigate this, covariate-adaptive randomization, which aims to balance treatment arms with respect to baseline prognostic factors, is commonly used. A popular method is stratified permuted block (SPB) randomization [Zelen, 1974]. However, practical constraints may prevent the inclusion of all baseline prognostic factors in covariate-adaptive randomization. In this simulation study, $X_{1}^{*}$ and $X_{2}^{*}$ are included in the SPB randomization, while all three covariates are used for adjustment during the analysis stage. In SPB randomization, subjects are categorized into strata based on $X_{1}^{*}$ and $X_{2}^{*}$, where $X_{2}^{*}$ is categorized into four quantiles. Within each stratum, subjects are further grouped into blocks of size 4, and permuted block randomization is applied within these blocks.
There are a total of 2,000 subjects. For each subject $i$, the death time $D_{i}=5/365+\omega_{i}$, where $\omega_{i}$ follows an exponential distribution with the mean as $\left\{0.05\hbox{exp}(X_{i}^{*T}\xi)\right\}^{-1}$, where $\xi=(0.1,0.1,0.1)^{T}$. The maximum follow-up time for each subject is 2 years. The independent censoring $C_{i}$ is uniformly distributed on $(1,2)$.
Recurrent event data are generated under the following five cases. Let $\theta$ denotes a scalar parameter controlling the magnitude of treatment effects, with $\eta=(0.2,0.2,0.2)^{T}$, and $\rho_{0}(t)=0.3t$. $\tau$ is set as 2 when calculating AUCs.
  • Case 1: Recurrent events are generated from the conditional intensity function that follows an Anderson-Gill model: $\rho^{(1)}_{j}(t\mid\boldsymbol{X}^{*})=\rho_{0}(t)\hbox{exp}(\theta j+\boldsymbol{X}^{*T}\eta)$ for $j=0,1$. The treatment effect characterized by $\theta$ remains constant over time.
  • Case 2: Recurrent events are generated using the conditional intensity function: $\rho^{(2)}_{j}(t\mid\boldsymbol{X}^{*})=\rho_{0}(t)\hbox{exp}\{\theta^{(1)}(t)j+\boldsymbol{X}^{*T}\eta\}$ for $j=0,1$, where $\theta^{(1)}(t)=-0.25\theta t^{2}+\theta t$. The treatment effect is characterized by a time-varying $\theta^{(1)}(t)$, which moves from 0 at $t=0$ to $\theta$ at $t=2$. This case features a late treatment effect scenario, where the treatment effect gradually increases and reaches its full extent by the end.
  • Case 3: Recurrent events are generated from the conditional intensity function: $\rho^{(3)}_{j}(t\mid\boldsymbol{X}^{*})=\rho_{0}(t)\hbox{exp}\{\theta^{(2)}(t)j+\boldsymbol{X}^{*T}\eta\}$ for $j=0,1$, where $\theta^{(2)}(t)=-\theta t^{2}+2\theta t$. $\theta(t)$ moves from 0 at $t=0$, reaches a peak of $\theta$ at $t=1$, and backs to 0 at $t=2$. This case exhibits a peak treatment effect scenario, with the effect being strongest at a specific point and weaker before and after that peak.
  • Case 4: Recurrent events are generated from the conditional intensity function: $\rho^{(4)}_{j}(t\mid\boldsymbol{X}^{*})=\rho_{0}(t)\hbox{exp}\{\theta^{(3)}(t)j+\boldsymbol{X}^{*T}\eta\}$ for $j=0,1$, where $\theta^{(3)}(t)=-0.25\theta t^{2}+\theta$. $\theta(t)$ moves from $\theta$ at $t=0$ to 0 at $t=2$. This case presents an early treatment effect scenario, where the effect is immediate at the beginning and then diminishes over time.
  • Case 5: Let $\Delta V_{ij}$ be the gap time between the $(j-1)$th and $j$th recurrent event for subject $i$. Define $\Delta V_{ij}=\hbox{exp}(-\theta I_{i}+X_{i}^{*T}\eta-0.7)+\epsilon_{j}$, where $\epsilon_{j}$ is an independent random variable following an exponential distribution with mean as 0.25. The ordered $j$th recurrent event time for subject $i$ is then $\sum_{k=1}^{j}\Delta V_{ik}$, for $\sum_{k=1}^{j}\Delta V_{ik}<T_{i}$, with the 00th recurrent event time being 0. Case 5 differs from the preceding cases in that the treatment affects recurrent events by reducing the gap time between them.
3.2 Results
The mean follow-up across the five cases and two randomization schemes is 1.44 years, with 7.8% of subjects experiencing mortality during the study period. The average number of recurrent events per subject is the same between the two randomization schemes but varies slightly between cases. For the placebo arm, the average number of recurrent events per subject is 0.68 for Cases 1 through 4 and 0.81 for Case 5. In the treatment arm, with $\theta=-0.16$, the corresponding averages are 0.58, 0.59, 0.61, 0.63, and 0.72 for Cases 1 through 5, respectively. When $\theta=-0.32$, the averages decrease further to 0.49, 0.51, 0.55, 0.58, and 0.62 for Cases 1 through 5, respectively.
Figure 1 shows the MCFs for the five cases with the simple randomization scheme, using a total sample size of 100,000 with equal allocation to generate smoother curves. $\theta$ is set to -0.32. The treatment effect is constant in Case 1 but not in Cases 2 to 5. However, it is difficult to detect the time-varying treatment effects from MCFs. To better view the treatment effect over time, the black curve in Figure 1 represents $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ over time. Despite some instabilities at earlier times, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ is constant over time in Case 1. In Case 2, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ decreases over time. In Case 3, the trend of $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ is not very clear before year 1, but after year 1, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ shows an obvious upward trend. In Case 4, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ increases over time, and in Case 5, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ remarkably increases early on and becomes stable near the tail.
Table 2 presents simulation results for the difference and ratio of AUCs from both unadjusted and covariate-adjusted analyses across various scenarios and randomization schemes, based on 5,000 simulations. For practical purposes, we use the Monte Carlo mean of the 5,000 estimated unadjusted differences and ratios of AUCs—denoted as $\Delta\hat{U}(\tau)$ and $\hat{R}(\tau)$—as empirical approximations of the true values $\Delta U(\tau)$ and $R(\tau)$, and label them as “Est.” “Bias” refers to the difference between the Monte Carlo mean of the 5,000 adjusted estimates ($\Delta\hat{U}^{(a)}(\tau)$ and $\hat{R}^{(a)}(\tau)$) and the corresponding “Est.” As expected, the “Bias” is consistently zero across all scenarios, empirically confirming that the proposed covariate adjustment procedure preserves the estimand. “Mean” and “Median” denote the mean and median of the 5,000 estimated standard errors (SEs) for each point estimate—$\Delta\hat{U}(\tau)$, $\Delta\hat{U}^{(a)}(\tau)$, $\hat{R}(\tau)$, and $\hat{R}^{(a)}(\tau)$—while “MC” represents the Monte Carlo standard deviation of the corresponding estimates. Coverage probability (CP) refers to the likelihood that the 95% confidence interval (CI) contains the respective “Est.” For the adjusted analyses, “Mean,” “Median,” and “MC” are nearly identical across all settings, and the CP remains close to the nominal level. In contrast, under the SPB randomization scheme, the unadjusted analysis yields “Mean” and “Median” values consistently larger than “MC,” with CP exceeding the nominal level—suggesting that unadjusted analyses that ignore covariate-adaptive randomization may be conservative. This observation aligns with FDA guidance on covariate adjustment, which recommends adjusting for baseline covariates in the analysis if they were balanced during randomization [Food and Drug Administration (2023), FDA]. Additional results for a smaller sample size (total of 400 subjects) are provided in Section D.1 of the Supplementary Materials. The results are similar to those presented in Table 2.
To visually demonstrate the validity of our variance formulas and highlight the conservativeness of the unadjusted analysis under covariate-adaptive randomization, Figure 2 displays the distribution of estimated standard errors $\sqrt{\hat{\sigma}_{L,R}^{2}/n}$ (unadjusted) and $\sqrt{\hat{\sigma}_{CL,R}^{2}/n}$ (adjusted) from 5,000 simulations under different randomization schemes and cases, with $\theta=-0.32$ and $\tau=2$. Similar results for $\sqrt{\hat{\sigma}_{L,\Delta}^{2}/n}$ and $\sqrt{\hat{\sigma}_{CL,\Delta}^{2}/n}$ are presented in Section D.2 of the Supplementary Materials. Yellow diamonds represent the Monte Carlo standard deviations of 5,000 $\hat{\mathcal{U}}_{L,R}$ (unadjusted) and $\hat{\mathcal{U}}_{CL,R}$ (adjusted). In the unadjusted analysis, the yellow diamond for simple randomization is higher compared to SPB randomization. This occurs because SPB randomization reduces variability by enforcing balance with respect to $X^{*}_{1}$ and $X^{*}_{2}$. For simple randomization in the unadjusted analysis, as well as for both simple and SPB randomization in the adjusted analysis, the boxplot bodies are thin and include the yellow diamond, despite some outliers. This empirically verifies the asymptotic properties outlined in Theorem 1, suggesting that our variance formulas are accurate and that the covariate adjustment approach is applicable to both simple and covariate-adaptive randomization. Conversely, in SPB randomization, the boxplots from the unadjusted analysis are positioned above the corresponding yellow diamonds, indicating that the estimated standard errors overestimate the actual standard deviations, making the analyses conservative. This observation further supports the notion that standard errors can be overestimated when covariate-adaptive randomization schemes are ignored.
Figure 3 compares the power of unadjusted AUC with adjusted AUC based on 5,000 simulations. Tests on the difference and ratio of AUCs show identical power. The figure demonstrates that the adjusted analysis consistently yields higher power across various cases and randomization schemes. This empirically supports the efficiency gain suggested by Theorem 1.
4 Real Examples
4.1 HF-ACTION
The HF-ACTION trial (Heart Failure: A Controlled Trial Investigating Outcomes of Exercise Training) was a multicenter, randomized controlled trial designed to evaluate the effects of structured exercise training on patients with chronic heart failure and reduced ejection fraction [O’Connor et al., 2009]. A total of 2,331 participants were randomly assigned to either an exercise training program or a control arm receiving usual care, stratified by center and heart failure etiology. The primary objective was to assess the impact of exercise training on a composite endpoint of all-cause hospitalization and mortality.
Our analysis focuses on a subgroup of 426 non-ischemic patients who had a baseline cardiopulmonary exercise test duration of nine minutes or less. Of these, 205 were assigned to the exercise training arm, and 221 to usual care. Over a median follow-up of approximately 28 months, mortality rates were about 18% in the exercise training arm compared to 26% in the usual care arm. The average number of recurrent hospitalizations per patient was 2.2 for the exercise arm and 2.6 for the usual care arm.
The upper panel of Figure 4 illustrates the mean cumulative function for both arms, along with the ratio of AUCs $\hat{U}_{1}(t)/\hat{U}_{0}(t)$ over time, represented by the black curve. Although there is some instability in the black curve at earlier time points, it remains relatively constant over time. This suggests that the constant treatment effect assumption of the LWYY model may not be significantly violated. However, for the purpose of illustrating our covariate adjustment approach for AUC, we will proceed with an AUC analysis.
In the analysis, we use an age indicator (whether a patient is over 60) as a covariate. Table 3 shows the estimated ratio and difference of AUCs between the two arms over the time window $[0,4]$. After adjusting for age, the variance of the log ratio of AUCs decreases from 0.0151 to 0.0147 (a 2.65% reduction), and the variance of the difference in AUCs decreases from 0.7695 to 0.7526 (a 1.02% reduction). These reductions suggest that covariate adjustment improves the precision of our estimates.
4.2 COPERNICUS
The COPERNICUS trial (Carvedilol Prospective Randomized Cumulative Survival) was a double-blind, placebo-controlled, multicenter study investigating the effect of carvedilol, a nonselective beta-blocker, on morbidity and mortality in patients with severe heart failure [Packer et al., 2001]. The study enrolled 2,289 participants, with 1,156 assigned to carvedilol and 1,133 to placebo. The primary outcome, death from any cause, occurred in 130 patients (11.2%) in the carvedilol arm compared to 190 patients (16.8%) in the placebo arm.
There were 700 hospitalizations in the carvedilol arm versus 848 in the placebo arm. The average number of recurrent hospitalizations per patient was 0.606 for the carvedilol arm compared to 0.748 for the placebo arm.
The lower panel of Figure 4 shows the MCF for both arms, along with the ratio of AUCs, $\hat{U}_{1}(t)/\hat{U}_{0}(t)$, over time (black curve). The treatment effect appears to remain relatively constant over time, making the LWYY model an appropriate choice for this data. However, for illustration purposes, we will proceed with an AUC analysis. Covariates in this analysis include sitting heart rate, history of myocardial infarction, presence of concomitant disease, diabetes, and the use of diuretics, ACE inhibitors, digitalis, amiodarone, and anticoagulants. Table 3 provides the estimated ratio and difference of AUCs between the two arms over the time window $[0,2.2]$. After adjusting for covariates, the variance of the log ratio of AUCs decreases from 0.0055 to 0.0052 (a 5.77% reduction), and the variance of the difference in AUCs decreases from 0.0153 to 0.0146 (a 4.79% reduction). These reductions further demonstrate that covariate adjustment improves the precision of our estimates.
4.3 Practical Implications
In this section, we demonstrate our covariate adjustment approach for AUC using two real-world clinical trials: the HF-ACTION trial and the COPERNICUS trial. For HF-ACTION, we have access only to a subgroup of participants with a single available covariate. The COPERNICUS trial includes all subjects, but some baseline covariates have missing values. While imputation is a viable strategy, we limit our analysis to covariates with complete data for illustrative purposes. Incorporating additional covariates would likely yield even greater efficiency gains.
The relative efficiency between two estimators is defined as the ratio of their variances. In clinical trial design, this can also be interpreted as the ratio of sample sizes required to achieve the same statistical power. In the COPERNICUS example, we observe an approximate 5% reduction in variance, which can be viewed as a potential 5% reduction in required sample size. A similar efficiency gain was observed in previous covariate adjustment studies [Ye et al., 2022, 2023, 2024].
Such efficiency gains are particularly meaningful in the context of modern clinical trials. As standard care continues to improve and event rates decline, larger sample sizes are often necessary to observe a sufficient number of events. For instance, while COPERNICUS (1997–2000) and HF-ACTION (2003–2007) enrolled fewer participants, many contemporary cardiovascular trials now exceed 10,000 subjects. A 5% reduction in required sample size could correspond to 500 fewer subjects in a 10,000-person trial—offering substantial savings in time, cost, and patient burden.
5 Discussion
Covariate-adaptive randomization is commonly used in clinical trials. Our simulation studies, along with previous research and FDA’s guidance, Food and Drug Administration (2023) [FDA] have demonstrated the necessity of adjusting for covariates balanced during randomization in subsequent analyses. Failure to do so can result in conservative estimates. In this paper, we aimed to improve the precision of AUC estimation with minimal assumptions under both simple and covariate-adaptive randomization, while preserving unconditional interpretations, using the nonparametric covariate adjustment approach proposed by Ye et al. [2022, 2023, 2024].
Theoretical asymptotic distributions demonstrate that our approach achieves efficiency gains, and extensive simulations across various scenarios and real data analysis confirm its effectiveness in improving statistical power. This nonparametric approach offers several advantages. First, this approach is model-free. Although we use simple linear regression to represent $\mathbb{E}(\hat{P}_{i}^{\Delta})$, $\mathbb{E}(\hat{Q}_{i}^{\Delta})$, $\mathbb{E}(\hat{P}_{i}^{R})$, and $\mathbb{E}(\hat{Q}_{i}^{R})$ during adjustment, the approach is valid regardless of whether the linear regression model correctly specifies the true expected values. In other words, its validity does not depend on parametric modeling assumptions [Zhang et al., 2008]. The choice of linear regression is for implementation convenience and offers good theoretical properties. This model-free characteristic aligns with the “minimal statistical assumption” principle advocated in the recent FDA guidance on covariate adjustment [Food and Drug Administration (2023), FDA]. Second, unlike conventional regression-based approaches for covariate adjustment, which can lose power if the model is misspecified, the nonparametric covariate adjustment guarantees an efficiency gain over the unadjusted test. Third, this approach is universally applicable to both simple and covariate-adaptive randomization schemes, unlike some nonparametric methods that are limited to simple randomization [Lu and Tsiatis, 2008]. Lastly, unlike non-linear regression models where the distinction between conditional and unconditional estimates is pronounced (known as non-collapsibility), this approach does not alter the unconditional interpretations.
In summary, three key principles underpin better practice in covariate adjustment: (1) preserving unconditional interpretation, (2) requiring minimal statistical assumptions, and (3) being universally applicable to both simple and covariate-adaptive randomization. Nonlinear regression models clearly do not satisfy principles (1) and (2). While some other covariate adjustment approaches satisfy principles (1) and (2), they are only applicable to simple randomization [Lu and Tsiatis, 2008]. The nonparametric covariate adjustment approach proposed by Ye et al. [2022, 2023, 2024] satisfies all three principles and is cited and recommended in the FDA guidance on covariate adjustment [Food and Drug Administration (2023), FDA].
While the unconditional treatment effect is often the primary focus in most randomized trials due to its relevance for broad policy recommendations and regulatory authorities [Tsiatis et al., 2008], some researchers argue that the conditional treatment effect is more appropriate for primary inferences. For instance, Hauck et al. [1998] advocate for adjusting for important prognostic covariates to achieve the most clinically relevant subject-specific measure of treatment effect. Inferences on the conditional effect can reveal interactions between treatment and patient characteristics, which may have significant implications for treatment use in specific subpopulations [Tsiatis et al., 2008]. Both unconditional and conditional treatment effects are important for a comprehensive understanding of treatment comparisons. However, this paper does not engage in this debate. Instead, we focus on covariate adjustment in the unconditional effect, aiming to improve the efficiency using minimal assumptions under different randomization schemes.
The AUC method has certain limitations. First, as discussed by Claggett et al. [2022], it is most appropriate when death rates are low and balanced across treatment arms. This condition is also most appropriate for the LWYY model. If a treatment prolongs survival, the AUC of the treatment arm is likely to be greater than that of the placebo arm, even if the recurrent event rate is lower in the treatment arm. One potential solution is to use the while-alive estimand [Schmidli et al., 2023, Wei et al., 2023, Mao, 2023]. Second, with small sample sizes, the MCF can be unstable near the tail, making the estimate sensitive to the selection of the time window for analysis. The time window should be prespecified in the study protocol but may be chosen empirically at analysis time with clinical justification [Tian et al., 2020].
To the best of our knowledge, this is the first paper to provide a robust, unconditional covariate-adjusted inference for the area under the curve estimand in the recurrent event setting, covering covariate-adaptive randomizations. The curve can represent not only the mean cumulative function but also the survival curve. The application of this approach to RMST is straightforward and discussed in Section E of the Supplementary Materials. While not the primary focus of this paper, we also examine the relative power of the LWYY and AUC approaches across various scenarios in Section D.3 of the Supplementary Materials, acknowledging that they test different null hypotheses. Future research on AUC may include sample size calculation and interim analysis.